​Colgate-Palmolive Company has just paid an annual dividend of $ 1.81. Analysts are predicting dividends to grow by $ 0.18 per year over the next five years. After​ then, Colgate's earnings are expected to grow 6.7 % per​ year, and its dividend payout rate will remain constant. If​ Colgate's equity cost of capital is 8.1 % per​ year, what price does the​ dividend-discount model predict Colgate stock should sell for​ today?

The price per share is $( ) (Round to four decimal​ places.)
Answer #1
Year Dividend
1 (1.81+0.18)=1.99
2 (1.99+0.18)=2.17
3 (2.17+0.18)=2.35
4 (2.35+0.18)=2.53
5 (2.53+0.18)=2.71

Value after year 5=(D5*Growth rate)/(Cost of capital-Growth rate)

=(2.71*1.067)/(0.081-0.067)

=$206.5407143

Hence current price=Future dividends and value*Present value of discounting factor(rate%,time period)

=1.99/1.081+2.17/1.081^2+2.35/1.081^3+2.53/1.081^4+2.71/1.081^5+$206.5407143/1.081^5

which is equal to

=$149.1660(Approx).